# Diffusion Posterior Sampling is Computationally Intractable

### v1

@@ Appendix A Lower Bound instance
For any $b\in\{-1,1\}$, let $\psi_{b}$ be the density of a standard Gaussian discretized to multiples of $\varepsilon$, with phase either 00 or $\frac{\varepsilon}{2}$ depending on $b$:
@@
See 3.1 We define our final Lower Bound distribution below, which is a scaled version of $g$. See 3.2
@@ B.1 Notation
Let $g$ be the distribution defined in 3.1, parameterized by a function $f$, and real values $R,\varepsilon>0$. For some noise pdf $h$, we define $\mathcal{X}^{h}_{f,R,\varepsilon}$ to be the distribution over $(x,y)$ where $x\sim g$ and $y\sim x_{[-d^{\prime}:]}+h$.
@@ B.2 Inverting $f$ via Posterior Sampling
[…]
@@
We know that $x^{b}$ is drawn from a uniform mixture over $g_{s}(x)$, as defined in 3.1. So, fixing an $s\in\{\pm 1\}^{d}$. We have that
@@
$\operatorname{\mathbb{P}}_{x^{b}}\left[\left|(x^{b}_{[:d]})_{i}-Rs_{i}\right|\leq\sqrt{2\log\frac{d}{\delta}}\right]\geq 1-\frac{\delta}{d}$
@@
$\operatorname{\mathbb{P}}_{x^{b}}\left[\text{Round}_{R}(x^{b}_{[:d]})=s\right]\geq 1-\delta.$ (4)
Putting together 2, 3, and 4 we get
$\operatorname{\mathbb{P}}_{x^{b}}\left[\text{Round}_{R}(x^{b}_{[:d]})=\text{Bits}_{\varepsilon}(y^{b})\right]\geq 1-\delta$
@@
[…]
@@
Let $\mathcal{X}^{\beta}_{f,R,\varepsilon}$ have pdf $p^{\beta}$. Assume we have a $(\tau,\delta)$-posterior sampler over $\mathcal{X}^{\beta}_{f,R,\varepsilon}$ that outputs sample from distribution $\widehat{\mathcal{X}}$ with distribution $\widehat{p}$. This means that with probability $1-\delta$ over $y$, there exists a coupling $\mathcal{P}$ over $(x,\widehat{x})$ such that $(x,\widehat{x})$ are $(\tau,\delta)$-close. Therefore, there exists a distribution $\mathcal{P}$ over $(x,\widehat{x},y)\in\mathbb{R}^{d+d^{\prime}}\times\mathbb{R}^{d+d^{\prime}}\times\mathbb{R}^{d^{\prime}}$ with density $p^{\mathcal{P}}$ such that $p^{\mathcal{P}}(x,y)=p^{\beta}(x,y)$, $p^{\mathcal{P}}(\widehat{x}\mid y)=\widehat{p}(\widehat{x}\mid y)$, and
$\operatorname{\mathbb{P}}_{x,\widehat{x}\sim\mathcal{P}}\left[\|x-\widehat{x}\|_{2}\leq\tau\right]\geq 1-2\delta.$
@@
$\operatorname{\mathbb{P}}_{x,\widehat{x}\sim\mathcal{P^{\prime}}}\left[\|x-\widehat{x}\|_{2}\leq\tau\right]\geq 1-2\delta$
@@
$\operatorname{\mathbb{P}}_{\widehat{x},x^{b}\sim\mathcal{P^{\prime}}}\left[\|\widehat{x}-x^{b}\|_{2}\leq\tau\right]\geq 1-3\delta$
@@
$\operatorname{\mathbb{P}}_{\widehat{x},x^{b}\sim\mathcal{P^{\prime}}}\left[\|\widehat{x}_{[:d]}-x^{b}_{[:d]}\|_{\infty}\leq\tau\right]\geq 1-3\delta.$ (5)
Now, by the definition of $\mathcal{X}^{\beta,\beta_{\text{max}}}_{f,R,\varepsilon}$, for all $i<d$, $x^{b}_{i}$ is a mixture of variance 1 normal distributions centered at ±$\pm R$. So, for any $i<d$,
$\operatorname{\mathbb{P}}_{x^{b}}\left[\left|x^{b}_{i}-\text{Round}_{R}(x^{b}_{i})\right|\geq\sqrt{2\log\frac{d}{\delta}}\right]\leq\frac{\delta}{d}$
Applying a union bound over $i\in[d]$ and putting this together with 5,
$\operatorname{\mathbb{P}}_{\widehat{x},x^{b}\sim P^{\prime}}\left[\|\widehat{x}_{[:d]}-\text{Round}_{R}(x^{b}_{[:d]})\|_{\infty}\leq\sqrt{2\log\frac{1}{\delta}}+\tau\right]\geq 1-4\delta$
@@
$\operatorname{\mathbb{P}}_{\widehat{x},x^{b}\sim P^{\prime}}\left[\|\text{Round}_{R}(\widehat{x}_{[:d]})-\text{Round}_{R}(x^{b}_{[:d]})\|_{\infty}\leq\sqrt{2\log\frac{d}{\delta}}+\tau\right]\leq 1-3\delta$
Again, the output of $\text{Round}_{R}$ is always ±$\pm R$, so this means
$\operatorname{\mathbb{P}}_{\widehat{x},x^{b}\sim P^{\prime}}\left[\text{Round}_{R}(\widehat{x}_{[:d]})=\text{Round}_{R}(x^{b}_{[:d]})\right]\geq 1-3\delta$
@@
$\operatorname{\mathbb{P}}_{x^{b},y^{b}\sim P^{\prime}}\left[f(\text{Round}_{R}(x^{b}_{[:d]}))=\text{Bits}_{\varepsilon}(y^{b})\right]\geq 1-\delta$
@@
$\operatorname{\mathbb{P}}_{\widehat{x},y^{b}\sim P^{\prime}}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))=\text{Bits}_{\varepsilon}(y^{b})\right]\geq 1-4\delta$
@@
$\operatorname{\mathbb{P}}_{\widehat{x},y\sim P^{\prime}}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))=\text{Bits}_{\varepsilon}(y)\right]\geq 1-5\delta$
@@
$\operatorname{\mathbb{P}}_{s,\mathsf{A}}[f(\mathsf{A}(f(s)))\neq f(s)]\leq 6\delta$
@@
$\operatorname{\mathbb{P}}\left[\text{Bits}_{\varepsilon}(y)=f(r)\right]\geq 1-\delta$
@@
$\operatorname{\mathbb{P}}_{\widehat{x},y}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))\neq\text{Bits}_{\varepsilon}(y)\right]\leq 5\delta$
@@
$\operatorname{\mathbb{P}}_{\widehat{x},y}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))\neq f(r)\right]\leq 6\delta$
@@ B.3 Inverting a One-Way function via Posterior Sampling
Since $d^{0.01}<m<d/2$, $d$ and $m$ are only polynomially separated. So, by G.2, we can construct a one-way function $f:\{\pm 1\}^{d-m}\to\{\pm 1\}^{m}$. By definition, we can see that $\widetilde{g}$, with measurement noise $\beta$ is the same distribution as $\mathcal{X}^{\beta R}_{f,R,\varepsilon}$, scaled down by $R$. Therefore, by Theorem B.4, since $R\geq 32\sqrt{\log\frac{d}{\delta}}$, $\varepsilon\geq\beta R\sqrt{\log\frac{d}{\delta}}$, if we can run a posterior sampler in time $T$, we can invert $f$ with probability $1-6\delta$ in time $T+O(m)$. So, if $f$ takes time superpolynomial in $m$ to invert, then $T+O(m)$ is superpolynomial. Since $m>d^{0.01}$, this means that $T$ itself is superpolynomial. ∎
@@
For any $m\leq d/2$, suppose that there exists a one-way function $f:\{\pm 1\}^{m}\to\{\pm 1\}^{m}$ that requires $2^{\Omega(m)}$ time to invert. Then, for $\widetilde{g}$ as defined in Definition 3.2 with $\varepsilon=\frac{1}{C\sqrt{\log d}}$ and $R=C\log d$, and linear measurement model with noise parameter $\beta=\frac{1}{C^{2}\log^{2}d}$ and measurement matrix $A\in\mathbb{R}^{m\times d}$, $(\frac{1}{10},\frac{1}{10})$-conditional sampling takes at least $2^{\Omega(m)}$ time.
@@
By definition, we can see that $\widetilde{g}$, with measurement noise $\beta$ is the same distribution as $\mathcal{X}^{\beta R}_{f,R,\varepsilon}$, scaled down by $R$. Therefore, by Theorem B.4, since $R\geq 32\sqrt{\log d}$, $\varepsilon\geq\beta R\sqrt{\log d}$, if we can run a posterior sampler in time $T$, we can invert $f$ with probability $0.4$ in time $T+O(m)$. So, if $f$ takes at least time $2^{\Omega(m)}$ to run, then we must have $T+O(m)\geq 2^{\Omega(m)}$, which means $T\geq 2^{\Omega(m)}$. ∎
@@ C.1 Piecewise Linear Approximation of $\sigma$-smoothed score in One Dimension
$\operatorname{\mathbb{P}}\left[\left|x-\mu\right|\geq\frac{m_{2}}{\delta}\right]\leq\delta$
@@
$=\sqrt{\operatorname*{\mathbb{E}}\left[s_{\sigma}(x)^{4}\right]\operatorname{\mathbb{P}}\left[\left|x-\mu\right|\geq\frac{m_{2}}{\sqrt{\delta}}\right]}$
@@
$=\frac{1}{\sigma^{2}}\log^{2}\frac{1}{\delta}\operatorname{\mathbb{P}}\left[|x-\mu|>\frac{m_{2}}{\sqrt{\delta}}\right]$
@@
$\leq\sqrt{\operatorname*{\mathbb{E}}\left[\left(\sup_{y\in[x-\gamma,x+\gamma]}s(x)\right)^{4}\right]\cdot\operatorname{\mathbb{P}}\left[m(x)>\frac{\log\frac{1}{\delta}}{\sigma}\right]}$
$\lesssim\sqrt{\frac{1}{\sigma^{4}}\cdot\operatorname{\mathbb{P}}\left[m(x)>\frac{\log\frac{1}{\delta}}{\sigma}\right]}$ by Lemma H.8
@@
As in Lemma C.1, this is the linear interpolation between $s(i\gamma)$ and $s((i+1)\gamma)$ on the interval $[i\gamma,(i+1)\gamma)$. By Lemma C.1, when $\gamma<\sigma/2$, we have
@@
$=\frac{\sqrt{\delta}}{\sigma^{2}}+\frac{1}{\sigma^{2}}\log^{2}\frac{1}{\delta}\operatorname{\mathbb{P}}\left[x\not\in G\right]$
$\leq\frac{\sqrt{\delta}}{\sigma^{2}}+\frac{1}{\sigma^{2}}\log^{2}\frac{1}{\delta}\operatorname{\mathbb{P}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(x)\geq\frac{1}{\sigma}\log\frac{1}{\delta}\right]$

### v2

@@ Appendix A Lower Bound instance
For any $b\in\{-1,1\}$, let $\psi_{b}$ be the density of a standard Gaussian discretized to multiples of $\varepsilon$, with phase either 0 or $\frac{\varepsilon}{2}$ depending on $b$:
@@
See 3.2 We define our final Lower Bound distribution below, which is a scaled version of $g$. See 3.3
@@ B.1 Notation
Let $g$ be the distribution defined in 3.2, parameterized by a function $f$, and real values $R,\varepsilon>0$. For some noise pdf $h$, we define $\mathcal{X}^{h}_{f,R,\varepsilon}$ to be the distribution over $(x,y)$ where $x\sim g$ and $y\sim x_{[-d^{\prime}:]}+h$.
@@ B.2 Inverting $f$ via Posterior Sampling
Let $\beta_{\text{max}}\leq\varepsilon/4$ and $\sqrt{32\log\frac{d}{\delta}}\leq R$. Then,
$\operatorname*{\mathbf{Pr}}_{x^{b},y^{b}\sim\mathcal{X}^{\beta,\beta_{\text{max}}}_{f,R,\varepsilon}}\left[f(\text{Round}_{R}(x^{b}_{[:d]}))=\text{Bits}_{\varepsilon}(y^{b})\right]\geq 1-\delta$
@@
We know that $x^{b}$ is drawn from a uniform mixture over $g_{s}(x)$, as defined in 3.2. So, fixing an $s\in\{\pm 1\}^{d}$. We have that
@@
$\operatorname*{\mathbf{Pr}}_{x^{b}}\left[\left|(x^{b}_{[:d]})_{i}-Rs_{i}\right|\leq\sqrt{2\log\frac{d}{\delta}}\right]\geq 1-\frac{\delta}{d}$
@@
$\operatorname*{\mathbf{Pr}}_{x^{b}}\left[\text{Round}_{R}(x^{b}_{[:d]})=s\right]\geq 1-\delta.$ (4)
Putting together eq. 2, eq. 3, and eq. 4 we get
$\operatorname*{\mathbf{Pr}}_{x^{b}}\left[\text{Round}_{R}(x^{b}_{[:d]})=\text{Bits}_{\varepsilon}(y^{b})\right]\geq 1-\delta$
@@
$\operatorname*{\mathbf{Pr}}[f(\text{Round}_{R}(\widehat{x}_{[:d]}))\neq\text{Bits}_{\varepsilon}(y)]\leq 5\delta.$
@@
Let $\mathcal{X}^{\beta}_{f,R,\varepsilon}$ have pdf $p^{\beta}$. Assume we have a $(\tau,\delta)$-posterior sampler over $\mathcal{X}^{\beta}_{f,R,\varepsilon}$ that outputs sample from distribution $\widehat{\mathcal{X}}$ with distribution $\widehat{p}$. This means that with probability $1-\delta$ over $y$, there exists a coupling $\mathcal{P}$ over (x,x^)($\widehat{x}$) such that (x,x^)($\widehat{x}$) are $(\tau,\delta)$-close. Therefore, there exists a distribution $\mathcal{P}$ over $(x,\widehat{x},y)\in\mathbb{R}^{d+d^{\prime}}\times\mathbb{R}^{d+d^{\prime}}\times\mathbb{R}^{d^{\prime}}$ with density $p^{\mathcal{P}}$ such that $p^{\mathcal{P}}(x,y)=p^{\beta}(x,y)$, $p^{\mathcal{P}}(\widehat{x}\mid y)=\widehat{p}(\widehat{x}\mid y)$, and
$\operatorname*{\mathbf{Pr}}_{x,\widehat{x}\sim\mathcal{P}}\left[\|x-\widehat{x}\|_{2}\leq\tau\right]\geq 1-2\delta.$
@@
$\operatorname*{\mathbf{Pr}}_{x,\widehat{x}\sim\mathcal{P^{\prime}}}\left[\|x-\widehat{x}\|_{2}\leq\tau\right]\geq 1-2\delta$
@@
$\operatorname*{\mathbf{Pr}}_{\widehat{x},x^{b}\sim\mathcal{P^{\prime}}}\left[\|\widehat{x}-x^{b}\|_{2}\leq\tau\right]\geq 1-3\delta$
@@
$\operatorname*{\mathbf{Pr}}_{\widehat{x},x^{b}\sim\mathcal{P^{\prime}}}\left[\|\widehat{x}_{[:d]}-x^{b}_{[:d]}\|_{\infty}\leq\tau\right]\geq 1-3\delta.$ (5)
Now, by the definition of $\mathcal{X}^{\beta,\beta_{\text{max}}}_{f,R,\varepsilon}$, for all $i<d$, $x^{b}_{i}$ is a mixture of variance 1 normal distributions centered at $\pm R$. So, for any $i<d$,
$\operatorname*{\mathbf{Pr}}_{x^{b}}\left[\left|x^{b}_{i}-\text{Round}_{R}(x^{b}_{i})\right|\geq\sqrt{2\log\frac{d}{\delta}}\right]\leq\frac{\delta}{d}$
Applying a union bound over $i\in[d]$ and putting this together with eq. 5,
$\operatorname*{\mathbf{Pr}}_{\widehat{x},x^{b}\sim P^{\prime}}\left[\|\widehat{x}_{[:d]}-\text{Round}_{R}(x^{b}_{[:d]})\|_{\infty}\leq\sqrt{2\log\frac{1}{\delta}}+\tau\right]\geq 1-4\delta$
@@
$\operatorname*{\mathbf{Pr}}_{\widehat{x},x^{b}\sim P^{\prime}}\left[\|\text{Round}_{R}(\widehat{x}_{[:d]})-\text{Round}_{R}(x^{b}_{[:d]})\|_{\infty}\leq\sqrt{2\log\frac{d}{\delta}}+\tau\right]\leq 1-3\delta$
Again, the output of $\text{Round}_{R}$ is always $\pm R$, so this means
$\operatorname*{\mathbf{Pr}}_{\widehat{x},x^{b}\sim P^{\prime}}\left[\text{Round}_{R}(\widehat{x}_{[:d]})=\text{Round}_{R}(x^{b}_{[:d]})\right]\geq 1-3\delta$
@@
$\operatorname*{\mathbf{Pr}}_{x^{b},y^{b}\sim P^{\prime}}\left[f(\text{Round}_{R}(x^{b}_{[:d]}))=\text{Bits}_{\varepsilon}(y^{b})\right]\geq 1-\delta$
@@
$\operatorname*{\mathbf{Pr}}_{\widehat{x},y^{b}\sim P^{\prime}}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))=\text{Bits}_{\varepsilon}(y^{b})\right]\geq 1-4\delta$
@@
$\operatorname*{\mathbf{Pr}}_{\widehat{x},y\sim P^{\prime}}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))=\text{Bits}_{\varepsilon}(y)\right]\geq 1-5\delta$
@@
$\operatorname*{\mathbf{Pr}}_{s,\mathsf{A}}[f(\mathsf{A}(f(s)))\neq f(s)]\leq 6\delta$
@@
$\operatorname*{\mathbf{Pr}}\left[\text{Bits}_{\varepsilon}(y)=f(r)\right]\geq 1-\delta$
@@
$\operatorname*{\mathbf{Pr}}_{\widehat{x},y}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))\neq\text{Bits}_{\varepsilon}(y)\right]\leq 5\delta$
@@
$\operatorname*{\mathbf{Pr}}_{\widehat{x},y}\left[f(\text{Round}_{R}(\widehat{x}_{[:d]}))\neq f(r)\right]\leq 6\delta$
@@ B.3 Inverting a One-Way function via Posterior Sampling
When $m>d/2$, we can add an arbitrary number of dummy observations which always observes 0. Posterior sampling in this instance is identical to only observing the first $d/2$ coordinates. Therefore, we only need to consider the case when $m\leq d/2$.
When $^{0.01}<m<d/2, dd$ and $m$ are only polynomially separated. So, by G.1, we can construct a one-way function $f:\{\pm 1\}^{d-m}\to\{\pm 1\}^{m}$. By definition, we can see that $\widetilde{g}$, with measurement noise $\beta$ is the same distribution as $\mathcal{X}^{\beta R}_{f,R,\varepsilon}$, scaled down by $R$. Therefore, by Theorem B.4, since $R\geq 32\sqrt{\log\frac{d}{\delta}}$, $\varepsilon\geq\beta R\sqrt{\log\frac{d}{\delta}}$, if we can run a posterior sampler in time $T$, we can invert $f$ with probability $1-6\delta$ in time $T+O(m)$. So, if $f$ takes time superpolynomial in $m$ to invert, then $T+O(m)$ is superpolynomial. Since $m>d^{0.01}$, this means that $T$ itself is superpolynomial in $d$. ∎
@@
Suppose that there exist one-way functions $f:\{\pm 1\}^{m}\to\{\pm 1\}^{m}$ that require $2^{\Omega(m)}$ time to invert. Then, for any $m=O(d)$, for $\widetilde{g}$ as defined in Definition 3.3 with $\varepsilon=\frac{1}{C\sqrt{\log d}}$ and $R=C\log d$, and linear measurement model with noise parameter $\beta=\frac{1}{C^{2}\log^{2}d}$ and measurement matrix $A\in\mathbb{R}^{m\times d}$, $(\frac{1}{10},\frac{1}{10})$-conditional sampling takes at least $2^{\Omega(m)}$ time.
@@
Similar to the proof of lemma 3.5, we only need to consider the case when $m\leq d/2$. By definition, we can see that $\widetilde{g}$, with measurement noise $\beta$ is the same distribution as $\mathcal{X}^{\beta R}_{f,R,\varepsilon}$, scaled down by $R$. Therefore, by Theorem B.4, since $R\geq 32\sqrt{\log d}$, $\varepsilon\geq\beta R\sqrt{\log d}$, if we can run a posterior sampler in time $T$, we can invert $f$ with probability $0.4$ in time $T+O(m)$. So, if $f$ takes at least time $2^{\Omega(m)}$ to run, then we must have $T+O(m)\geq 2^{\Omega(m)}$, which means $T\geq 2^{\Omega(m)}$. ∎
@@ C.1 Piecewise Linear Approximation of $\sigma$-smoothed score in One Dimension
$\operatorname*{\mathbf{Pr}}\left[\left|x-\mu\right|\geq\frac{m_{2}}{\delta}\right]\leq\delta$
@@
$=\sqrt{\operatorname*{\mathbb{E}}\left[s_{\sigma}(x)^{4}\right]\operatorname*{\mathbf{Pr}}\left[\left|x-\mu\right|\geq\frac{m_{2}}{\sqrt{\delta}}\right]}$
@@
$=\frac{1}{\sigma^{2}}\log^{2}\frac{1}{\delta}\operatorname*{\mathbf{Pr}}\left[|x-\mu|>\frac{m_{2}}{\sqrt{\delta}}\right]$
@@
$\leq\sqrt{\operatorname*{\mathbb{E}}\left[\left(\sup_{y\in[x-\gamma,x+\gamma]}s(x)\right)^{4}\right]\cdot\operatorname*{\mathbf{Pr}}\left[m(x)>\frac{\log\frac{1}{\delta}}{\sigma}\right]}$
$\lesssim\sqrt{\frac{1}{\sigma^{4}}\cdot\operatorname*{\mathbf{Pr}}\left[m(x)>\frac{\log\frac{1}{\delta}}{\sigma}\right]}$ by Lemma H.8
@@
As in Lemma C.1, this is the linear interpolation between $s(i\gamma)$ and $s((i+1)\gamma)$ on the interval $i\gamma,(i+1)\gamma)$. By Lemma C.1, when $\gamma<\sigma/2$, we have
@@
$=\frac{\sqrt{\delta}}{\sigma^{2}}+\frac{1}{\sigma^{2}}\log^{2}\frac{1}{\delta}\operatorname*{\mathbf{Pr}}\left[x\not\in G\right]$
$\leq\frac{\sqrt{\delta}}{\sigma^{2}}+\frac{1}{\sigma^{2}}\log^{2}\frac{1}{\delta}\operatorname*{\mathbf{Pr}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(x)\geq\frac{1}{\sigma}\log\frac{1}{\delta}\right]$